OpenStudy (pulsified333):

Find the equation of the tangent line to the graph of f(x)=2xln(x+7) at x=−5.

OpenStudy (jdoe0001):

well, we know x = -5 and y = f(x) and so, the equation of the tangent line will be \(\bf y-f(x_0)=m(x-x_0)\implies y-f(x_0)={\color{brown}{ \frac{dy}{dx}}}(x-x_0)\)

OpenStudy (jdoe0001):

so, "x" is given, you can get "y" from f(x) all you need really is the slope, or derivative, at x = -5 and plug it in the point-slope form :)

OpenStudy (jdoe0001):

keep in mind that \(\bf f(x)=2xln(x+7)\implies f(x)=ln\left[ (x+7)^{{\color{brown}{ 2x}}} \right] \)

OpenStudy (pulsified333):

okay